Evelyn and Eleanor worked a total of 36 hours. If Evelyn worked 10 hours longer than Eleanor, how many hours did each woman work?
Answer by jim_thompson5910(35256)

Let x = Evelyn's hours and y = Eleanor's hours

"Evelyn and Eleanor worked a total of 36 hours" ---> x+y = 36

"Evelyn worked 10 hours longer than Eleanor" ---> x = y+10
-------------------------------------------------------
x+y = 36

(y+10)+y = 36

y+10+y = 36

2y+10 = 36

2y = 36-10

2y = 26

y = 26/2

y = 13

------------

x = y+10

x = 13+10

x = 23

So Evelyn worked 23 hours and Eleanor worked 13 hours.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------